# Spectrum of slip dynamics, scaling & statistical laws emerge from simplified model of fault and damage zone architecture

### v1

@@ 3.1 Seismic Cycles and Slip Dynamics
In Figure 4b, one slow slip event is showcased, rupturing only part of the main fault. This slow slip event spans 279 days with a moment magnitude $M_{w}=4.48$. It is essential to acknowledge the subjectivity in detecting slow events due to the slip rate threshold used (here defined by slip rates between $10^{-9}$ and $10^{-3}$ m/s). This threshold effect will be discussed later in the text. We observe continuous seismic activity on the off-fault fractures accompanying the slow event on the main fault, as previously seen in Figure 3c. Figure 4c unfolds a full rupture on the main fault with a moment magnitude of $M_{w}=5.03$, revealing nuanced variations in slip rate acceleration and deceleration within the rupture and along the fault’s strike. The nucleation and afterslip of this event are 110 and 12 days, respectively. Figure 4d features another full rupture on the main fault ($M_{w}=4.95$), preceded by a nucleation phase that ruptures around one-third of the main fault over 58 days. The rupture accelerates from the left corner of the nucleation patch. Then, the full rupture on the main fault lasts for 39 seconds; after a 45-day afterslip phase, the slip rate falls below $10^{-9}$ m/s. The figure illustrates a distinct cluster of off-fault events following the rupture front on the main fault.
4 Properties of the events in the catalog
@@
In this section, we present the results of our numerical models, demonstrating their ability to replicate observed empirical statistical relationships and scaling laws governing seismic activity. These empirical relationships include the Omori law for the decay in the rate of aftershocks, Gutenberg-Richter magnitude-frequency distribution, the inverse-Omori law for foreshock escalation, moment-area scaling, and moment-duration scaling. We posit that any reasonable ‘digital twin’ of a fault system should be able to reproduce these empirical laws in addition to the range of slip dynamics. Through this analysis, we plan to show that our fault volume model not only produces a diverse range of slip dynamics but also satisfies key empirical laws observed in natural seismicity.
@@ 4.1 Omori Law
In seismological studies, the Omori law characterizes the temporal decay of aftershock activity following a mainshock event. This empirical law, commonly known as the Omori law, describes how the rate of aftershocks decreases over time following a seismic event. Specifically, it states that the number of aftershocks decreases inversely with time after the mainshock, following a power-law decay as follows: $n(t)=k/(c+t)^{p}$ where $p$ is a constant that describes the decay rate and typically falls in the range 0.8–1.2 in most cases [139, 197]. In Figure 5a, we visualize the aftershock activity decay using stacked sequences from multiple seismic cycles within a simulation. Each gray line represents the aftershock activity following a single major seismic event, while the blue line represents the stacked sequences, providing an aggregated view of the aftershock decay pattern. By fitting the stacked sequences to the Omori law, we find a compelling match with an exponent $p$ of 0.92, indicating a consistent decay rate across the simulated seismic events.
Figure 5: (a) Omori Law: Seismicity rate decreases over time following the mainshock. Each gray curve represents one earthquake cycle with one mainshock on the main fault, the blue curve represents the stacked sequences of all the seismic cycles in this simulation case (b) Magnitude-frequency distribution of the cataloged fast ruptures follows the Gutenberg-Richter distribution. Red and blue colors represent fast and slow ruptures respectively, while black color represents the total catalog (c) Inverse Omori: Seismicity rate increases inversely with time as the main rupture approaches. Each gray curve represents one earthquake cycle with one mainshock on the main fault, the blue curve represents the stacked sequences of all the seismic cycles in this simulation case (d) Scaling laws of inferred seismicity: Moment area scaling. Events are color-coded based on rupture velocity $v_{r}$.
4.2 Gutenberg-Richter Magnitude Frequency Distributions
@@
Figure 5b illustrates the magnitude-frequency distribution of ruptures from one of the simulations in our study. The x-axis denotes the magnitude range, while the y-axis represents the frequency of occurrences. Different distributions are plotted for fast, slow, and all ruptures. The distribution follows the Gutenberg-Richter law, exhibiting a logarithmic relationship between magnitude and frequency. This distribution provides crucial insights into the relative occurrence rates of seismic events of varying magnitudes, offering valuable information for seismic hazard assessment and earthquake forecasting. We can observe that the fast ruptures dominate the catalog in this case. To estimate the b-value, we utilize the maximum likelihood method [3], compensating for the binning error [118]. We note that $M_{w}=-2$ approximately serves as the completeness magnitude of our catalog. Our measured b-value for fast ruptures, approximately 0.5. As shown by [4], a b-value between 0.5 and 1 can be envisaged when the fractal dimension is between 1 and 2 where the fault lines are distributed along a plane. This needs to be examined in much more detail. It is important to note that our primary objective is not to study the origin of the b-value derived from our fault volume model. Rather, our focus is to demonstrate that considering the dynamics of a fault volume with fractures surrounding the main fault results in a logarithmic relationship between magnitude and frequency, consistent with the observed Gutenberg-Richter law in nature. While we suspect that the b-value may be influenced by factors such as the power-law distribution of fracture length and density distribution of fractures along fault-normal distance, we defer this exploration to future studies, as it falls beyond the scope of our current investigation.
@@ 4.3 Inverse Omori Law
In addition to the Omori-law-like reduction in the earthquake rate after a given event (section 4.1), the catalog of events in this paper’s simulation follows the empirical inverse Omori law that characterizes the increase in foreshock activity preceding a major rupture event. Unlike the aftershock decay described by the traditional Omori law, the inverse Omori law predicts a gradual rise in foreshock activity leading up to a significant seismic event, where foreshock rate increases as an inverse of the time to the mainshock as follows: $n(t)=1/(c+\Delta t)^{p}$ [91, 181], where $p$ is a constant that describes the increase rate of foreshocks and is typically around 1. In Figure 5c, we present the seismicity rate leading up to rupture events on the main fault, with each gray line representing the foreshock activity during a single seismic cycle. The blue line denotes the stacked sequences, providing an aggregated view of the precursor activity preceding significant ruptures. By fitting the stacked sequences to the inverse Omori law, we obtain a notable alignment with a parameter value of $p\approx 0.74$, close to 1. The alignment with the Omori law underscores its applicability in describing the temporal evolution of aftershock activity within our simulation. Similarly, the consistent trend of foreshock rate increase observed across our simulated seismic events aligns well with the predictive power of the inverse Omori law, demonstrating its ability to capture precursory behavior preceding major seismic events and laboratory earthquakes as shown in [117].
4.4 Scaling Laws
@@
Next, in Figure 5d, we explore the relationship between seismic moment ($M_{0}$) and rupture area ($S$) for both fast and slow ruptures. Notably, while determining rupture area ($S$) in nature can often be challenging, our model offers a straightforward approach as we have precise knowledge of fault behavior. Given our 2D model, we assume a circular rupture shape, simplifying the calculation of rupture area as equal to the square of rupture length. The figure reveals a scaling relation of $M_{0}\propto S^{3/2}$, mirroring patterns observed in natural seismicity [95]. It is essential to note that this scaling law is consistent across both fast and slow ruptures.
Figure 6: Scaling laws of inferred seismicity: Moment-duration scaling with a threshold of detection of slow ruptures at (a) $10^{-6}$ m/s and (b) $10^{-8}$ m/s. Events are color-coded based on rupture velocity $v_{r}$.
@@
In Figure 6a, we delve into the moment-duration scaling, a crucial aspect of seismic rupture behavior. Observations in natural seismicity have highlighted a cubic relation between moment and duration for fast ruptures: $M_{0}\propto T^{3}$ [94, 96, 80]. However, for slow ruptures, the situation appears more complicated, with some observations suggesting a linear relation: $M_{0}\propto T$ [82, 81], while others indicating a cubic one [70, 122]. To investigate this further, we assess whether the detection threshold for slow ruptures influences the observed scaling relationship. Specifically, we examine Figure 6b, where the detection threshold is set to $10^{-8}$ m/s, and compare it to Figure 6a, which uses a threshold of $10^{-6}$ m/s. This comparison allows us to evaluate how the choice of detection threshold affects the apparent moment–duration scaling. Notably, the latter threshold is still three orders of magnitude smaller than the threshold for fast ruptures, set at $10^{-3}$ m/s. The color scale represents the logarithm of rupture velocity. The bottom x-axis displays seismic moment ($M_{0}$), while the top x-axis shows moment magnitude ($M_{w}$) for clarity and comparison with observational data. For the fast ruptures, we observe a clear cubic scaling ($M_{0}\propto T^{3}$) between moment and duration, consistent with previous findings. However, the behavior of slow ruptures varies between the two figures. In Figure 6b with the lower detection threshold, we observe a linear relation between moment and duration. In contrast, Figure 6a with the higher threshold shows a trend closer to cubic scaling. This scaling is more apparent when we compile the entire catalog of all the simulations that we ran, and is displayed later in Figure 14. The effect of the detection threshold, ranging from $10^{-9}$ m/s to $10^{-6}$ m/s, on the detectable moment and magnitude of an example slow-slip rupture is shown in Figure 7a-d, highlighting how inferences drawn from the duration and moment of a slow-slip rupture could be changed if the lowest slip rate at which the rupture is initiated cannot be detected. This discrepancy suggests that the scaling between moment and duration for slow ruptures is highly dependent on the detection threshold in our numerical model, akin to the sensitivity or threshold settings of field recording instruments or detection strategies employed on field data. In a recent work [47], similar threshold-dependent effects were observed, in natural settings, when applying deep learning denoising techniques to reveal a continuum of slow slip events in the Cascadia subduction zone, demonstrating that detection methodology fundamentally shapes our understanding of slip behavior. Although the slip rate on the main fault may be very low and remain undetectable, the pattern of off-fault seismicity migration provides valuable information about the rupture process on the main fault, Figure 7e. The dot symbols on the plot represent off-fault events projected onto the main fault, marking the extents of the slow slip rupture. Further analysis is warranted to assess how reliably the pattern of off-fault seismic activity can be used to infer the magnitude and duration of a slow event, and we defer this investigation to future studies.
@@
Since each of the events is color-coded based on its average rupture velocity in Figure 6, it is apparent that slow or fast events are identical in nature except for their rupture velocity. In fact, it is evident that $M_{0}\propto T^{3}$ and $M_{0}\propto T$ are limiting bounds in the moment-duration scaling as proposed by [81]. This suggests that slow ruptures are simply fast ruptures with relatively lower slip rates, rupture velocities, and stress drops. This also implies that the slow ruptures do not emerge due to a special frictional consitutive law since it is the same mechanism that generates fast ruptures as well. What thus produces the difference in rupture velocity is the spatiotemporally complex traction on the main fault due to the presence of the fault volume.

### v2

@@ 3.1 Seismic Cycles and Slip Dynamics
In Figure 3b, one slow slip event is showcased, rupturing only part of the main fault. This slow slip event spans 279 days with a moment magnitude $M_{w}=4.48$. It is essential to acknowledge the subjectivity in detecting slow events due to the slip rate threshold used (here defined by slip rates between $10^{-9}$ and $10^{-3}$ m/s). This threshold effect will be discussed later in the text. We observe continuous seismic activity on the off-fault fractures accompanying the slow event on the main fault, as previously seen in Figure 2c. Figure 3c unfolds a full rupture on the main fault with a moment magnitude of $M_{w}=5.03$, revealing nuanced variations in slip rate acceleration and deceleration within the rupture and along the fault’s strike. The nucleation and afterslip phases of this event last 110 and 12 days, respectively. Figure 3d features another full rupture on the main fault ($M_{w}=4.95$), preceded by a nucleation phase that ruptures around one-third of the main fault over 58 days. The rupture accelerates from the left corner of the nucleation patch. Then, the full rupture on the main fault lasts for 39 seconds; after a 45-day afterslip phase, the slip rate falls below $10^{-9}$ m/s. The figure illustrates a distinct cluster of off-fault events following the rupture front on the main fault.
Beyond demonstrating that slow and fast events emerge naturally within the same fault system under identical constitutive parameters, our model is also able to produce spatially localized regions where slow slip is persistent. This emerges naturally from the local geometric configuration of the main fault and surrounding off-fault fractures, without requiring imposed frictional or hydraulic heterogeneity. The coexistence of slow and fast slip events on the same fault or within the same region (see Figure S2), as produced by our simulations, is also supported by a growing body of observational, geological, and laboratory evidence. Seismological observations document the coexistence or close spatial association of slow slip, tremor, very-low-frequency earthquakes, and regular earthquakes in several tectonic settings [[, e.g.]]Kato2012, Johnson2012, Ito2013, Ruiz2014, Tsang2015, Lin2020, Villafuerte2025, Woods2024. Geological studies indicate that deformation mechanisms associated with slow and fast slip can operate within the same fault zones over geological timescales [[, e.g.]]Fagereng2019, and laboratory experiments demonstrate the coexistence and interaction of slow and fast slip under controlled conditions [[, e.g.]]Aben2023.
3.2 Slow to fast events across scales
Figure 4: Moment-rate functions (MRFs) (left panels), and their spectra (right panels), across different $M_{w}$ ranges for both off-fault and main fault events. (a) Selected MRFs from slow and fast events in the off-fault region within a narrow range of $0.9\leq M_{w}\leq-0.8$. (b) Same as (a), but for $0.3\leq M_{w}\leq-0.2$. (c) MRFs from main fault events, and their spectra, with fast ruptures in the range $2.2\leq M_{w}\leq 3.8$ and slow slip events in the range $2.4\leq M_{w}\leq 4.8$.
We present a general overview of the source time functions (STFs) generated within our fault system model, focusing on their variability in moment magnitude, duration, and spectral characteristics. The analysis includes moment-rate functions (MRFs) from both slow and fast events, considering a representative fault volume geometry in which off-fault fractures are oriented parallel to the main fault (unlike the case discussed earlier where the off-faults are optimally oriented right-lateral faults).
While we get a continuum of slow and fast events for all off-fault geometries, we selected this particular geometry because it exhibits a remarkably broad and diverse range of temporal behaviors, capturing both small-magnitude slow events and large slow slip events occurring on both off-faults and the main fault. This suggests that adjusting frictional parameters could potentially produce similar behavior across different geometries, though this requires further investigation.
Figure 5: MRF shapes for slow and fast events after normalization by centroid time with colors indicating moment magnitude. Panels (a) and (b) correspond to slow events in the off-fault region for the same magnitude ranges as in Figures 4a and 4b. (c) shows slow events on the main fault. Panels (d) and (e) correspond to fast off-fault events with the same $M_{w}$ ranges as in (a) and (b). (f) Fast events on the main fault.
Figures 4a and 4b show selected MRFs from off-fault events within narrow magnitude ranges: $0.9\leq M_{w}\leq-0.8$ and $0.3\leq M_{w}\leq-0.2$, respectively. These MRFs exhibit a continuous and broad distribution of slip durations, with significant variability in their shape, ranging from abrupt moment release to smoother and more gradual evolutions. While the fast events are classical earthquakes, the slower events can easily be called LFE’s and VLFEs. To confirm this, we need to radiate these events dynamically and look at their far-field waveforms, which we plan to do in future work. Although not explored here, our model framework could also explain tectonic tremors as successive ruptures of shear cracks as recently proposed by [192]. This diversity highlights the influence of structural complexity and distributed deformation in controlling rupture behavior, emphasizing that seismic complexity arises not only from the main fault, as shown previously, but also from off-fault activity.
Figure 4c shows MRFs of events on the main fault, which exhibit a bimodal behavior. Due to the greater extent and roughness, the main fault hosts the largest magnitude events, characterized by more complex STFs, including multiple acceleration and deceleration phases during the ruptures. The magnitude range considered here includes both large-magnitude, fast ruptures, partial or full, ($2.2\leq M_{w}\leq 3.8$) and long-duration slow slip events ($2.4\leq M_{w}\leq 4.8$).
We now examine the spectral characteristics of these events. The corresponding source spectra for the same events show, as expected, that the longer durations of the slow events lead to significantly lower corner frequencies than fast events across all three magnitude ranges, on both the main and off-faults. Recently, [189] analyzed the S-wave displacement spectral signature of low-frequency earthquakes in the Nankai Trough, after correcting for empirically derived attenuation, and concluded that the spectra are consistent with the classical earthquake model with vastly different rupture velocities and stress drops. Our results align very well with the conclusions drawn by [189]. Beyond the corner frequency, the spectral decay of slow events appears similar to that of fast events. However, the high-frequency content is less reliably estimated, as it would be predominantly influenced by the fully dynamic response of both event types, an aspect that our quasi-dynamic model does not account for, but which we aim to address in future investigations. To further examine MRF characteristics across these events, we normalized the time axis of each MRF with respect to its centroid time, $T_{c}$ (Figure 5). The centroid time is the moment rate weighted average time [55, 54] and is defined as $T_{c}={\int_{t_{1}}^{t_{2}}t\dot{M}_{0}(t)\,dt}/{\int_{t_{1}}^{t_{2}}\dot{M}_{0}(t)}$. Here $\dot{M}_{0}(t)$ is the moment rate, and $t_{1}$ and $t_{2}$ correspond to the onset and termination times of the MRF, defined by when the slip rate crosses a detection threshold.
One key distinction between slow and fast STFs lies in their symmetry. MRFs of slow events (Figures 5a–c) tend to be more asymmetric, typically exhibiting a sharp rise followed by a longer, gradual decay—indicating a positively skewed moment release. In contrast, fast earthquakes generally show more symmetric MRFs with a single, well-defined peak (Figures 5d–f). However, for the largest slow slip events (Figure 5c), the STFs of both fast and slow events become more irregular, and the asymmetry for some events is less distinguishable. Consistent with the findings of [112], our results suggest that fast earthquakes tend to have simpler MRFs, with fewer subevents and a well-defined peak and gradual decay (Figures 5d–f). While this section focuses on a specific fault geometry, similar results were obtained for all the 40 fault volume configurations tested in this study.
4 Statistical and Scaling Laws of the Events in the Catalog
@@
In this section, we demonstrate the ability of our numerical model to replicate observed empirical statistical relationships and scaling laws governing seismic activity. These empirical relationships include the Omori law for the decay in the rate of aftershocks, Gutenberg-Richter magnitude-frequency distribution, the inverse-Omori law for foreshock escalation, moment-area scaling, and moment-duration scaling. We posit that any reasonable ‘digital twin’ of a fault system should be able to reproduce these empirical laws in addition to the range of slip dynamics presented earlier. Through this analysis, we plan to show that our fault volume model not only produces a diverse range of slip dynamics but also satisfies key empirical laws observed in natural seismicity. We show that for each geometrical configuration and friction parameters, the above listed empirical laws are satisfied (see Figures 2e, S3, S4). For the sake of brevity here we show the robustness of these results across all the different cases of off-fault orientations and frictional parameters.
@@ 4.1 Omori Law
In seismological studies, the Omori law characterizes the temporal decay of aftershock activity following a mainshock event. Specifically, it states that the number of aftershocks decreases inversely with time after the mainshock, following a power-law decay as follows: $n(t)=k/(c+t)^{p}$ where $k$ is aftershock productivity, $c$ is a time offset, and $p$ is a constant that describes the decay rate and typically falls in the range 0.8–1.2 in most cases [130, 184]. To evaluate the Omori law in our simulations, for each orientation distribution of the off-fault fractures, we stack the aftershock sequences following each major rupture on the main fault. We then fit the stacked sequences to the Omori law to estimate the $p$ exponent. Figure 6a illustrates the Omori law coefficient $p$ across different fracture orientations when $a/b=0.5$. The coefficient remains consistently around 1 across all orientations, indicating a uniform behavior regardless of these parameters. While off-fault orientation does influence the duration of the aftershock period (see the case when off-faults are aligned with the main fault), we defer this investigation to future studies as it falls outside the scope of our current research.
Figure 6: (a) Omori Law: Effect of off-fault fracture orientation. The average Omori decay curve over a period of 30 days is calculated from the stacked curves of different fracture orientation, and the average $p$ coefficient is reported (b) Magnitude Distribution: Effect of off-fault fracture orientation on the b-value found from stacked catalog of fast ruptures with same off-fault fracture orientation (c) Inverse Omori: Effect of off-fault fracture orientation (d) Moment-area scaling: The $a/b$ ratio of the off-fault fractures is 0.5.
4.2 Gutenberg-Richter Magnitude Frequency Distribution
@@
Figure 6b illustrates the magnitude-frequency distribution of ruptures across different fracture orientations. The x-axis denotes the magnitude range, while the y-axis represents the frequency of occurrences. The distribution follows the Gutenberg-Richter law, exhibiting a logarithmic relationship between magnitude and frequency. This distribution provides crucial insights into the relative occurrence rates of seismic events of varying magnitudes, offering valuable information for seismic hazard assessment and earthquake forecasting. To estimate the b-value, we utilize the maximum likelihood method [3], compensating for the binning error [111]. We note that $M_{w}=-2$ approximately serves as the completeness magnitude of our catalog. All plotted distributions conform to the Gutenberg-Richter law, displaying a similar $b$-value across orientations. Interestingly, there is no apparent influence of off-fault fracture orientation on the magnitude-frequency distribution, suggesting that this aspect does not significantly impact slip dynamics.
As shown by [4], a b-value between 0.5 and 1 can be envisaged when the fractal dimension is between 1 and 2 where the fault lines are distributed along a plane. This needs to be examined in much more detail. Although there is still room to investigate the convexity of the magnitude–frequency distribution using truncated G–R distributions [72], we leave this investigation for future studies. It is important to note that our primary objective is not to study the origin of the b-value derived from our fault volume model. Rather, our focus is to demonstrate that considering the dynamics of a fault volume with fractures surrounding the main fault results in a logarithmic relationship between magnitude and frequency, consistent with the observed Gutenberg-Richter law in nature. While we suspect that the b-value may be influenced by factors such as the power-law distribution of fracture length and density distribution of fractures along fault-normal distance, we defer this exploration to future studies, as it falls beyond the scope of our current investigation.
@@ 4.3 Inverse Omori Law
In addition to the Omori-law-like reduction in the earthquake rate after a given event (section 4.1), the catalog of events in this paper’s simulation follows the empirical inverse Omori law that characterizes the increase in foreshock activity preceding a major rupture event. Unlike the aftershock decay described by the traditional Omori law, the inverse Omori law predicts a gradual rise in foreshock activity leading up to a significant seismic event, where foreshock rate increases as an inverse of the time to the mainshock as follows: $n(t)=1/(c+\Delta t)^{p}$ [84, 169], where $p$ is a constant that describes the increase rate of foreshocks and is typically around 1. In Figure 6c, we show the increasing rate of seismicity in a 3-hour window prior to the main shock, grouped by fracture orientation. The increasing rate of seismic events off-fault exhibits an inverse Omori law, with $p$-values close to 1. The alignment with the Omori law underscores its applicability in describing the temporal evolution of aftershock activity within our simulation. Similarly, the consistent trend of foreshock rate increase observed across our simulated seismic events aligns well with the predictive power of the inverse Omori law, demonstrating its ability to capture precursory behavior preceding major seismic events and laboratory earthquakes as shown in [110].
4.4 Moment-Area Scaling Law
@@
Next, in Figure 6d, we explore the relationship between seismic moment ($M_{0}$) and rupture area ($S$) for both fast and slow ruptures. Notably, while determining rupture area ($S$) in nature can often be challenging, our model offers a straightforward approach as we have precise knowledge of fault behavior. Given our 2D model, we assume a circular rupture shape, simplifying the calculation of rupture area as equal to the square of rupture length. The figure reveals a scaling relation of $M_{0}\propto S^{3/2}$, mirroring patterns observed in natural seismicity [88]. It is essential to note that this scaling law is consistent across both fast and slow ruptures. This scaling law underscores the fault volume model’s fidelity in capturing the spatial distribution of seismic energy release across various rupture sizes. Importantly, the observation that both fast and slow ruptures exhibit almost the same scaling between moment and rupture area is noteworthy. Given that seismic moment relates to stress drop and rupture area as $M_{0}=C\Delta\sigma S^{3/2}$ [56, 88, 106], where $C$ is a geometric constant ($C=16/7\pi^{3/2}$ for a circular crack), the variability in stress drop must be smaller than that of rupture velocity. In fact the above relationship can be rearranged to show that $\overline{V}\mu/\overline{v_{r}}\Delta\sigma$ is a constant. Here $\overline{V}$ is the average slip rate of the event and $\overline{v_{r}}$ is the average rupture velocity of the event. In our simulations, average sliprates vary over 8 to 9 orders of magnitude between slow and fast events, while rupture velocities vary over 5 to 6 orders of magnitude. This implies that stress drops should vary over 2 to 3 orders of magnitude across the entire catalog of events. It is also clear that average slip rate, stress drop and average rupture velocity are three fundamental quantities of an event that cannot vary independently and are linked through the above relationship. [178], for instance, found that low-frequency earthquakes propagate more slowly and have lower stress drops than regular earthquakes.
4.5 Moment-Duration Scaling Law
@@
In Figure 7, we delve into the moment-duration scaling, a crucial aspect of slow and fast ruptures. Observations in natural seismicity have highlighted a cubic relation between moment and duration for fast ruptures: $M_{0}\propto T^{3}$ [87, 89, 71]. However, for slow ruptures, the situation appears more complicated, with some observations suggesting a linear relation: $M_{0}\propto T$ [74, 75], while others indicating a cubic one [62, 115]. To investigate this further, we assess whether the detection threshold for slow ruptures influences the observed scaling relationship. Specifically, we examine Figure 7b, where the detection threshold is set to $10^{-9}$ m/s, and compare it to Figure 7a, which uses a threshold of $10^{-6}$ m/s. This comparison allows us to evaluate how the choice of detection threshold affects the apparent moment–duration scaling. Notably, the latter threshold is still three orders of magnitude smaller than the threshold for fast ruptures, set at $10^{-3}$ m/s. The color scale represents the rupture velocity. The bottom x-axis displays seismic moment ($M_{0}$), while the top x-axis shows moment magnitude ($M_{w}$) for clarity and comparison with observational data. For the fast ruptures, we observe a clear cubic scaling ($M_{0}\propto T^{3}$) between moment and duration, consistent with previous findings.
@@
However, the behavior of slow ruptures varies between the two detection thresholds. With a lower threshold (Figure 7b), the durations of slow events are larger and we also obtain more events compared to the higher threshold case (Figure 7a). The overall trend of the slow events thus changes as a function of the detection threshold. This discrepancy suggests that the scaling between moment and duration for slow ruptures is highly dependent on the detection threshold in our numerical model, akin to the sensitivity or threshold settings of field recording instruments or detection strategies employed on field data. In a recent work [42], similar threshold-dependent effects were observed, in natural settings, when applying deep learning denoising techniques to reveal a continuum of slow slip events in the Cascadia subduction zone, demonstrating that detection threshold fundamentally shapes our understanding of slip behavior. The effect of the detection threshold, ranging from $10^{-9}$ m/s to $10^{-6}$ m/s, on the detectable moment and magnitude of an example slow-slip rupture is shown in Figure 8a-d, highlighting how inferences drawn from the duration and moment of a slow-slip rupture could be changed if the lowest slip rate at which the rupture is initiated cannot be detected.
While the slip rate on the main fault may be very low and remain undetectable, the pattern of off-fault seismicity migration provides valuable information about the rupture process on the main fault, Figure 8e. The dot symbols on the plot represent off-fault events projected onto the main fault, marking the extents of the slow slip rupture. Further analysis is warranted to assess how reliably the pattern of off-fault seismic activity can be used to infer the magnitude and duration of a slow event, and we defer this investigation to future studies.
Since each of the events is color-coded based on its average rupture velocity in Figure 7, it is apparent that slow or fast events are identical in nature except for their rupture velocity. In fact, it is evident that $M_{0}\propto T^{3}$ and $M_{0}\propto T$ are limiting bounds in the moment-duration scaling as proposed by [75]. This suggests that slow ruptures are simply fast ruptures with different slip rates, rupture velocities, and stress drops. This also implies that the slow ruptures do not emerge due to a special frictional constitutive law since it is the same mechanism that generates fast ruptures as well. What thus produces the difference in rupture velocity is the spatiotemporally complex traction on the main fault due to the presence of the fault volume. We also observe small events tend to have a continuum of rupture velocities; however, as events become larger, a distinct gap emerges between fast and slow ruptures. This gap mirrors observations reported by [75], suggesting events that are challenging to detect or exceedingly rare. Our numerical model’s ability to detect these events with confidence reinforces the presence of this gap, raising the possibility of a mechanical constraint where large events predominantly exhibit either fast or slow rupture velocities, rather than a continuum. This strongly suggests that both fast and slow ruptures do originate from the same mechanical model, despite their differing rupture velocities.